OpenStudy (samigupta8):

all the value of the parameter a for which inequality given below is satisfied by atleast one real x is A.a>=2 B.a>2 C.a<=2

OpenStudy (samigupta8):

\[4^{x}-a2^{x}-a+3\le 0\]

OpenStudy (samigupta8):

i solve it using hit and trial bt i m little confused about a=2 and answer is first option

hartnn (hartnn):

let 2^x be y then 4^x = y^2 got this ?

OpenStudy (samigupta8):

yaa

OpenStudy (samigupta8):

the equation wud be y^2-ay-a+3<=0

OpenStudy (samigupta8):

wt happnd is is it wrong

hartnn (hartnn):

y^2-ay +(3-a) <=0 Ay^2+By+C <=0 so find A,B,C from here, then for solutions to be real B^2-4AC >=0

OpenStudy (samigupta8):

bt there is a+6 also wht to do vid dat

OpenStudy (samigupta8):

answer wud be -infinity to 6 also

OpenStudy (samigupta8):

@hartnn pls...reply